how do I find the equation of a line that passes through 1,2 and is perpendicular to y equals negative three fourths X plus two and three fourths
Answer by mananth(16946)

1 y = - 3/ 4 x + 11/4
Divide by 1
y = - 3/ 4 x + 11/4
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 3/4

The slope of a line perpendicular to the above line will be the negative reciprocal 4/3
Because m1*m2 =-1
The slope of the required line will be4/3

m=4/3 ,point ( 1 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = 4/ 3 + b
b= 2/3
m= 4/3
The required equation is y = 4/ 3 x + 2/ 3
m.ananth@hotmail.ca